OpenStudy (anonymous):

solve x^2 - 8= -16 by completing the square

hero (hero):

Hint add (-8/2)^2 to both sides

OpenStudy (anonymous):

is it 7,8?

hero (hero):

x^2 - 8x + 16 = -16 + 16 x^2 - 8x + 16 = 0 (x - 4)^2 = 0 x - 4 = 0 x = ?

OpenStudy (anonymous):

4,4

hero (hero):

x = 4 There's only one solution for x. You do not write 4 twice in this case.

OpenStudy (anonymous):

okay. thats my only option on the answer choices